A beam of light in air strikes a plate of crown glass at an angle of incidence of
40°. What is the angle of refraction?
1
Expert's answer
2020-11-16T07:37:36-0500
"n_1\sin{\theta_1}=n_2\sin{\theta_2}\\\theta_2=\arcsin{\frac{\sin{40}}{1.52}}=25\degree"